OpenStudy (anonymous):

f(x)= 7^x logbase3 (x) I need to find the derivative .... but I can only enter the answer using change of base - can anyone provide direction? Thanks

OpenStudy (anonymous):

Is it correct

OpenStudy (anonymous):

\[7^x \log_3(x)\]

OpenStudy (anonymous):

\[d/dx \log(x)=1/xln(b)\]

OpenStudy (anonymous):

scratch that

OpenStudy (anonymous):

\[d/dx \log_b(x)=1/xln(b)\]

OpenStudy (anonymous):

So you will use product rule

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

ok I see that and I am kinda clear - in that eq. where is the 7^x ah..1/7ln3

OpenStudy (anonymous):

3ln(x)=1/7ln3

OpenStudy (anonymous):

How is going so far?

OpenStudy (anonymous):

is it...? [(1\7ln 3(x)) -(3\ln (x))]

OpenStudy (anonymous):

\[d/dx (7^x)=\ln (7)*7^x\]

OpenStudy (anonymous):

product rule is u*v'+v*u'

OpenStudy (anonymous):

my computer died. I am back and will try again. Thank you now I'm guest.... this is 1477

OpenStudy (anonymous):

haven't quite got there but I have to go now. thank you so much for your help!!!